### v1

# Honeybee: Decentralized Peer Sampling with Verifiable Random Walks for Blockchain Data Sharding

Abstract.
Data sharding—in which block data is sharded without sharding compute—is at the present the favored approach for scaling Ethereum. A key challenge toward implementing data sharding is verifying whether the entirety of a block’s data is available in the network (across its shards). A central technique proposed to conduct this verification uses erasure coded blocks and is called data availability sampling (DAS). While the high-level protocol details of DAS has been well discussed in the community, discussions around how such a protocol will be implemented at the peer-to-peer layer are lacking. We identify random sampling of nodes as a fundamental primitive necessary to carry out DAS and present Honeybee, a decentralized algorithm for sampling node that uses verifiable random walks. Honeybee is secure against attacks even in the presence of a large number of Byzantine nodes (e.g., 50% of the network). We evaluate Honeybee through experiments and show that the quality of sampling achieved by Honeybee is significantly better compared to the state-of-the-art. Our proposed algorithm has implications for DAS functions in both full nodes and light nodes.
blockchain, peer-to-peer, data availability sampling
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Symposium on Principles of Distributed Computing; June 17–21, 2024; Nantes, Franceprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Networks Network algorithmsccs: Theory of computation Distributed algorithmsccs: Networks Peer-to-peer protocolsccs: Networks Network performance evaluationccs: Networks Network propertiesccs: Networks Peer-to-peer networksccs: Computing methodologies Multi-agent systems
1. Introduction
Blockchains are steadily maturing into ecosystems that support decentralized applications (dapp) in diverse domains—including finance, payments, storage, games, healthcare etc.—and used by millions of clients conducting billions of dollars of transactions each day (CoinMarketCap, 2024). As demand for dapp usage increases, it is important that the blockchains can handle the high rate of transaction requests. Today the Ethereum blockchain can process at most a few ten transactions per second on average which has resulted in unacceptably high transaction fees during periods of high demand (Etherscan, 2024).
To address the scaling problem, the Ethereum community is focused on the development of a data-sharding design in which (1) block sizes are increased from the current 80 kB (average) to as big as 30 MB (Nikolaenko and Boneh, 2024) and erasure coded, (2) each validator node stores only a small chunk of each coded block. An increased block size naturally admits a greater number of transactions and improves transaction throughput of the chain. Combined with the use of layer-2 scaling methods, particularly Rollups (Ethereum.org, 2024e), Ethereum envisions a roadmap wherein a theoretical maximum throughput of 100,000 transactions per second is feasible (Forum, 2024). Data sharding in Ethereum is colloquially referred as Dank sharding, named after its proposer (EIP-4844, 2024).
A fundamental challenge toward realizing data sharding is the design of the peer-to-peer (p2p) network. Unlike Ethereum 1.0 which used a simple broadcast gossip primitive to disseminate transactions and blocks in the network or Ethereum 2.0 which uses a publish-subscribe model to disseminate different types of messages (e.g., attestations, blocks, transactions) over different subnets, a node in a data-sharded Ethereum must additionally be able to check the availability of a full block by sampling random chunks from other nodes (Al-Bassam et al., 2018). This requires a requesting node to be able to contact randomly selected nodes in the network and download chunks. Having the ability to sample random chunks is important not only for full functional nodes, but also for light nodes (e.g., a wallet running on a smart phone) which can number in the millions compared to the thousands of full nodes available today (Ethernodes.org, 2024).
In an open, permissionless and decentralized setting that Ethereum operates on, developing an algorithm by which a node can uniformly sample other nodes in the p2p network is far from straightforward. At present Ethereum’s p2p stack (based on devp2p and libp2p) use the Kademlia distributed hash table (DHT) protocol (Maymounkov and Mazieres, 2002) to sample and discover new nodes, as part of the node discovery procedure (to-peer Networking Specifications, 2024b, a). To discover a random node in the network, all a node has to do is issue a query for a randomly selected target identifier in Kademlia and receive the IP address information of the node that is closest to the target identifier (in Kademlia’s XOR distance sense).
Even though Ethereum uses Kademlia as the protocol of choice to perform network-level node sampling, in this paper we argue that in the presence of adversarial nodes (e.g., Sybils), Kademlia does not achieve uniform sampling and can cause a requesting node to become eclipsed. As a potential replacement to Kademlia, in this paper we propose Honeybee which is a fully decentralized p2p algorithm for performing uniform node sampling even in the presence of adversarial nodes. A Honeybee node achieves sampling by participating in several random walks over the p2p overlay. Each node in Honeybee maintains an address table containing addresses of the most recently sampled peers which are also used to progress random walks of other peers visiting the node. To protect against adversarial attacks, Honeybee uses verifiable randomness derived from the blockchain to perform the walks. Additionally, Honeybee nodes also perform peer-to-peer reconciliation of node address tables to identify and expose attackers engaging in equivocating their address tables.
Sampling nodes in a distributed system using random walks has historically been well-studied in the context of applications such as overlay monitoring, design of expanders, search, routing, resource management etc (Awan et al., 2006). However, prior works in this space consider models that do not simultaneously satisfy our requirements: (1) the algorithm must be decentralized, (2) there can be a large number (e.g., constant fraction) of adversarial nodes, and (3) adversarial nodes can exhibit arbitrary Byzantine behavior (e.g., message insertions, deletions during gossip). E.g., Anceaume et al. (Anceaume et al., 2013a, b) consider achieving sampling through streaming messages between neighbors, but assume there are no Sybil attacks. Augustine et al. (Augustine et al., 2015) consider a dynamic p2p network model where the attacker decides how the network churns from round to round. However, within each round the assumption is that gossiping happens without any message loss. The works Awan et al. (Awan et al., 2006), Gkantsidis et al. (Gkantsidis et al., 2004), propose distributed algorithm for sampling in unstructured p2p networks, but do not consider adversarial node behavior.
We do not attempt to present a full-fledged p2p network design for data sharding in Ethereum in this paper. Rather we posit that a uniform node sampling capability will have a central role to play in the (eventual) overall network design. The sampling capability can also enhance the effectiveness of broadcast in today’s Ethereum p2p subnets. [^1] In the context of Ethereum’s publish-subscribe p2p network, a subnet is an overlay wherein all nodes are interested in the same topic and gossip messages on that topic. The emerging trend of modular blockchains advocates for a separate data availability layer in the blockchain stack, the implementation of which can make use of Honeybee as well (Celestia, 2024; Docs, 2024a).
We evaluate Honeybee through a custom simulator we have built. [^2] Code will be made open source. Compared to the baseline algorithms GossipSub and Kademlia, Honeybee achieves the same level of near-uniform sampling when all nodes in the network adhere to the protocols. However, when the network contains adversarial nodes, Honeybee outperforms GossipSub and Kademlia by 4-63% in terms of sampling adversarial peer ratio. We define that an algorithm achieves $\epsilon$-uniform sampling when the sampling adversarial peer ratio from the algorithm is bounded from above by the sum of the true adversarial nodes ratio in the network and $\epsilon$. Under such standard, Honeybee consistently achieves 0.03-uniform sampling with 5-50% adversarial nodes in the network.
2. Background
2.1. Data availability sampling
Transactions per second, or throughput, is a key measure of blockchain scalability. Today the Ethereum blockchain has an average block size of less than 200 kB corresponding to a throughput of around 30 transactions per second. Increasing this throughput to 1000s of transactions per second (to the scale of Visa, for example) has been a long-standing open challenge in the community.
While several layer-2 solutions such as state channels, side channels, Plasma etc. (Ethereum.org, 2024b) have been proposed to improve throughput, a particularly important layer-2 solution that has become the focus of Ethereum’s scaling roadmap is the rollup (Ethereum.org, 2024e). Briefly, a rollup is an independent blockchain that is bridged to the Ethereum’s main chain via a smart contract. For increased security in the rollup chain, blocks produced in the rollup are periodically published on the main chain in a compressed form. This allows any verifier on the main chain to verify the rollup operation, and slash rollups block producers in case of a mistake. The rollups blocks published on Ethereum are stored as data ‘blobs’ and are not executed by Ethereum validators by default.
Including blobs from rollups within an Ethereum block can significantly amplify the size of the block to several megabytes. In this scenario, requiring all the nodes to store all the blocks (as is the case today) can overwhelm resource-limited nodes and affect the decentralization of the network. Therefore, Ethereum envisions a sharded design in which a block is split into many smaller chunks, and each node is required to store only a small number of the chunks (Ethereum.org, 2024a; Al-Bassam et al., 2018). The block header is stored by all the nodes.
The challenge is to ensure that a published block can be successfully reconstructed from its chunks later if required. E.g., a malicious block publisher may reveal only the block header, retaining the block body; or the publisher may publish only 90% of the chunks; or some of chunks can be modified during publication etc. To solve this, a block is first erasure coded to increase its size by 4 times (Ethereum uses a 2D-Reed Solomon code). The erasure coded block is then divided into chunks which are then dispersed over the validator network.
Erasure coding the blocks is advantageous in two ways:
  1. (1)
    It is possible for a node to reconstruct the entire block by fetching any random 25% of the chunks. This is easier to achieve compared to retrieving all 100% of the chunks if a block is not erasure coded.
  2. (2)
    It is simple for a node to verify that a block’s data is available (i.e., has been dispersed correctly). All the node needs to do is sample a small $k$ number of random chunks from the network. If the node can successfully retrieve all $k$ chunks, then with probability $>1-1/4^{k}$, block data is available over the network. This is because if a malicious node attempts to withhold publishing of a portion of the block, then it can do it only by withholding more than 75% of the coded chunks.
Correctness of each individual chunk is verified using polynomial commitments (specifically, KZG commitment) that are packaged along with the chunks. The process of checking whether adequate block data is present in the network by sampling is called data availability sampling.
Data availability sampling is essential not only for full nodes but is also beneficial to light nodes (Ethereum.org, 2024c). Light nodes are end users interacting with the blockchain through lightweight application software than can run easily on resource-limited devices. A simple example is a user running a wallet software on her browser that interacts with the blockchain through the wallet provider’s servers. As illustrated by this example, today light nodes must trust a centralized server (typically a full node) for submitting and confirming transactions, and checking validity of blocks.
As the Ethereum network scales, there is an increasing need support light nodes without requiring trust on a central party. Efforts such as the Portal network (Docs, 2024b) hint at the community’s desire to progress in this direction. We envision a design in which light nodes connect with other light nodes to form a gossip network. Full nodes supply block headers to light clients which gossip the header to other clients over the gossip network. In case a block is invalid, full nodes construct a fraud proof which is again gossiped over the light node network. However, a full node can construct a fraud proof only if all the data in the invalid block is available in the network. Thus, it is important for light nodes to independently verify data availability of a block lest they think the block is valid sans fraud proof.
2.2. Ethereum Peer-to-Peer Network
Ethereum uses a p2p publish-subscribe (pub-sub) network based on the GossipSub protocol (Vyzovitis et al., 2020) from the libp2p framework (Libp2p, 2024) for disseminating messages. In a pub-sub network, each node subscribes to one or more topics the node is interested in. Any message published by a node in the network belongs to a unique topic. The network is set up such that when a message is published from a topic, all nodes subscribed to that topic receive the message. In GossipSub this is achieved by constructing one independent overlay for each topic, containing all nodes subscribed to that topic. Therefore, whenever a message is published from a topic, by simply broadcasting the message over the topic’s overlay, all subscribers are guaranteed to receive the message. The primary use of a pub-sub network in Ethereum is for aggregating attestations from 64 different subsets of validators (of Stake Consensus Specifications, 2024).
In the context of data sharding, once again Ethereum plans to use a pub-sub network for data distribution and retrieval. Specifically, if a block is divided into $m$ chunks (after erasure coding), there will be $m$ topics with the $i$-th topic including the $i$-th chunk of each block. Each node subscribes to a small number of topics and is responsible for storing chunks of those topics. Chunk dissemination is straightforward: the block publisher disseminates the $i$-th chunk of the block over the $i$-th subnet. Details on how a node looking to randomly sample chunks can contact nodes from a different topic overlay to fetch chunks are unfortunately lacking in publicly available documents, forums and blog-posts.
3. System Model
3.1. Network and Security Model
We consider a network comprising of $n$ nodes $V$, out of which a fraction $f$ of the nodes are adversarial. Nodes that are not adversarial are called honest. We denote the set of honest and adversarial nodes by $V_{h}$ and $V_{a}$ respectively. Each node has a unique network address (i.e., IP address, port). A node can connect and communicate with another node if it knows the latter’s network address. Each node has a small memory of size $k$, for storing information about $k$ other nodes in the network. [^3] We refer to the memory space also as an address table. Apart from the network address, we assume a node also has a public, private key pair which can be used for signing messages, issuing commitments etc. Since the memory space is small, an honest node cannot know the network addresses of the entire network. However, adversarial nodes can pool together the addresses they know of to conduct attacks.
We consider Byzantine adversaries in that they can arbitrarily deviate from our proposed protocol. Examples of Byzantine behavior include not responding to or arbitrarily delaying requests or sending malicious messages to victims. We assume the network connection between nodes is reliable (i.e., a synchronous model), and do not model message loss or delays.
When a node first joins the network, it contacts a bootstrapping server from which it receives information (network address, public key) about $k$ random nodes in the network. It is common for practical p2p networks (including Ethereum’s) to use bootstrapping servers with hardcoded IP address to help new peers join the network (Ethereum.org, 2024d). Once a node joins the network, it must run its own discovery protocol and cannot query the bootstrapping server for fresh addresses. Note that the initial set of addresses the bootstrapping server provide can include adversarial nodes as well.
We assume the public key to network address binding of a node is attested by the bootstrapping server through a certificate signed by the server. While it is possible to use a decentralized public-key infrastructure for this purpose (Shi et al., 2022), we consider the bootstrapping server as the trusted certificate authority in our setting for simplicity.
Time is divided into rounds $t=0,1,2,\ldots$. During a round, a node can send or receive messages of total size $l$, where $l$ is again a small constant that models the bandwidth constraints of the node. Lastly, we assume nodes have access to a fresh public random number each round. In practice, a new block is produced in Ethereum every 12 seconds. Thinking of 12 seconds as a round, a random number can be derived each round from the header of the block for that round. Note that we require nodes to only download the block header to compute the randomness, and not the entire block, which is particularly useful for light nodes.
3.2. Problem Statement
For any node $v\in V$, let $M_{v}(t)$ denote the contents of the address table (i.e., memory) of node $v$ at time $t$. For any $i\in\{1,2,\ldots,k\}$, let $M_{v}^{i}(t)$ be the $i$-th address in $M_{v}(t)$. Our objective is to design a decentralized algorithm $\pi$ by which an honest node can uniformly sample nodes in the network. Specifically, for any node $v\in V_{h}$, time $t$ and index $i\in\{1,2,\ldots,k\}$, we want
(1) $P(M_{v}^{i}(t)\in V_{h})$ $\geq 1-f$
(2) $P(M_{v}^{i}(t)=u)$ $=P(M_{v}^{i}(t)=u’)$ ’ )
for any honest nodes $u,u’\in V_{h}$. The reason we lower bound the probability is achieving perfect uniform sampling would be impossible if the adversarial nodes do not participate in the protocol. On the other hand if adversarial nodes all behave honestly perfectly uniform random sampling must be possible.
The requirements outlined above alone can be trivially satisfied if a node downloads random addresses from the bootstrapping server and does nothing afterwards. Therefore, we qualify our objective by additionally requiring that a node must add at least one fresh sample to its address table every $\Delta>0$ rounds, i.e.,
(3) $M_{v}(t+\Delta)\backslash M_{v}(t)\neq\{\},$
for all $t>0$. Furthermore, for any time $t$ let $R_{v}(t)$ be the most recently added address to $v$’s address table. We want the newly sampled node to be independent of the past samples, i.e.,
(4) $P(R_{v}(t)=u|M_{v}(0),M_{v}(1),\ldots,M_{v}(t-\Delta))=P(R_{v}(t)=u),$
for all $u\in V_{h}$ and $t>\Delta$.
4. Honeybee
4.1. Algorithm Overview
Honeybee  is a fully decentralized p2p algorithm that tackles the data availability sampling problem on the network level by conducting secure near-uniform [^4] Note that near-uniform is well-defined for peer sampling. sampling of peers from the entire network. Honeybee  is inspired by random walks on finite groups and mixing Markov chains (Aldous, 1983; Lovász, 1993). A Honeybee  node maintains three tables: an address table, an encounter table, and a connection table. In the address table, a Honeybee  node stores peer data for peers with whom the node has peering agreements. The address table consists of two sub-tables: an outgoing address table that contains data for at most $n_{out}$ peers and an incoming address table that contains data for at most $n_{in}$ peers. If a node $A$ sends a request to peer with node $B$, and node $B$ accepts the request, node $A$ will add node $B$ to $A$’s outgoing address table, while node $B$ will add node $A$ to $B$’s incoming address table. The peering agreement in Honeybee  is bilateral - node $A$ having node $B$ in $A$’s address table would suggest that node $B$ having node $A$ in $B$’s address table. To be specific, when a node $A$ has node $B$ in its address table, node $A$ stores the following data in its address table: $B$’s node ID, $B$’s address table snapshot, $B$’s IP address, $B$’s port information, and a peering agreement signed by both $A$ and $B$ with a timestamp. A node’s address table snapshot contains everything in its address table except for its peers’ address table snapshots. A node’s encounter table is a finite FIFO list of size $n_{e}$ that contains the address table snapshots of a random subset of the nodes it has interacted with via random walks. A node’s connection table contains the peers with whom the node communicate.
(a) Overview of Honeybee VRW
(b) Illustration of Honeybee VRW
(c) Illustration of Honeybee TCC
Figure 1. Honeybee illustrations: (a) shows an example of a Honeybee  verifiable random walk sampling, and node $C$ replaces node $B$ in node $A$’s outgoing address table; (b) shows how a verifiable random walk (VRW) works; (c) shows how a table consistency check (TCC) works.
To join the Honeybee  network, a node first add some of the bootstrap nodes for an initial configuration. The bootstrap nodes are a small group of trustworthy nodes that not only participate in the network but also assist other nodes with configurations. In Honeybee, all nodes follow the blockchain for the shared pseudo-randomness (i.e., block header), and the shared pseudo-randomness determines the nodes that are eligible to perform sampling at the current moment (currently eligible nodes). To sample peers from the network in a near-uniform way, a currently eligible node conducts a random walk with the path determined by the shared pseudo-randomness. A random walk message carries its initiator node information, destination node information, and the random walk path information. During a random walk, the node on each hop of the random walk conduct table consistency checks with the initiator node. When a random walk message reaches its destination node, the message serves as a request for the random walk initiator node to peer with the destination node. If the destination node accepts the request, the initiator node and the destination node add each other to their address tables and establish a peering agreement. For a node, we call the time between successive roles as a currently eligible node as an epoch (round). In Algorithm 1, we present the overall template for outgoing address table update of Honeybee. We give the definitions of verifiable random walks and table consistency checks in §4.2 and §4.3. In Fig. 1a, we give an example of a Honeybee  verifiable random walk sampling. Fig. 1b shows how a verifiable random walk works. Fig. 1c shows how a table consistency check works.
input : peers $\Gamma^{v}_{\text{curr}}$ in outgoing address table of current epoch; pseudo-randomness seed $\mathcal{R}$ indicating whether node $v$ is a currently eligible node; $v$’s secret key $SK_{v}$; epoch counter $\varrho_{\text{curr}}$ enumerating the number of epochs since $v$ joins the network; outgoing address table size $n_{out}$; system-defined address table inconsistency threshold $\tau$; bootstrap node addresses $b$;
output : updated set of peers $\Gamma^{v}_{\text{next}}$ for next epoch; updated epoch counter $\varrho_{\text{next}}$ for next epoch;
/* Perform a random walk if $v$ is a currently eligible node */
if $\mathcal{R}$ indicates $v$ is a currently eligible node then
       $p\leftarrow\textsc{GetPathLength}(\mathcal{R},v,SK_{v})$ $p_{c}\leftarrow 0$ /* Hop counter of the random walk */
       $u^{*}\leftarrow v$ /* Node on current hop of the random walk */
       $d\leftarrow\emptyset$ /* $d$ stores the node on the first hop of the random walk */
       while $p_{c}<p$ do
             /* VRF() returns a pseudo-random number $i$ and a proof $\pi_{v}$/
             $(i,\pi_{v})\leftarrow VRF(\mathcal{R},\varrho_{\text{curr}},p_{c},u^{*},SK_{v})$ /* $\Gamma^{u^{*}}_{\mathrm{curr}}(i)$ returns i-th node in $\Gamma^{u^{*}}_{\mathrm{curr}}$ and its address table snapshot */
             $(u^{*},\omega)\leftarrow\Gamma^{u^{*}}_{\mathrm{curr}}(i)$ $p_{c}\leftarrow p_{c}+1$ if $p_{c}=1$ then
                   $d\leftarrow u^{*}$
             end if
            $\omega^{*}\leftarrow\textsc{RequestAddrTable}(u^{*})$ if $\textsc{Diff}(\omega,\omega^{*})>\tau$ then
                   $\textsc{IssueFraudProof}(u^{*},\omega,\omega^{*})$
             end if
       end while
      /* GetDestResponse() returns the response $\varphi$ of the destination node */
       $\varphi\leftarrow\textsc{GetDestResponse}(u^{*})$ if $\varphi$ indicates $u^{*}$ accepts request from $v$ then
             $\Gamma^{v}_{\text{next}}\leftarrow\Gamma^{v}_{\text{curr}}\backslash\{d\}\cup\{u^{*}\}$
       end if
      /* Add new peers if outgoing address table is not full */
       if Size$(\Gamma^{v}_{\text{curr}})$ $<$ $\mathrm{n_{out}}$  then
             $\Gamma^{v}_{\text{next}}\leftarrow\Gamma^{v}_{\text{curr}}\cup\textsc{RequestNewPeers}(b,n_{out}-\textsc{Size}(\Gamma^{v}_{\text{curr}}))$
       end if
      $\varrho_{\text{next}}\leftarrow\varrho_{\text{curr}}+1$
end if
Algorithm 1 Algorithm outline for updating entries of outgoing address table of node $v$ in each epoch.
4.2. Verifiable Random Walks
A Honeybee  node performs a random walk to replace one of its outgoing table peers when it learns that it becomes a currently eligible node from the pseudo-randomness seed (i.e., block header). However, a random walk is susceptible to adversarial routing. If a random walk from a honest node meets a dishonest node, the dishonest node can route the random walk to another dishonest node and eventually guide the random walk to a dishonest destination. To ensure our random walks are truly random, we employ a verification mechanism in Honeybee’s random walks (VRW). When a node $A$ learns that it is a currently eligible node, it derives a verifiable pseudo-random number and a proof $\pi_{A}$ with its secret key using the following inputs: the current block header, its epoch count, its random walk’s current hop count, and the public key of the node on its current hop. The generated verifiable pseudo-random number informs node $A$ where to proceed with the next hop of its random walk. Along with the signature of the node on the current hop, the nodes on the next hop of the random walk can verify the validity of the random walk. The procedure is repeated for each hop until the end of the random walk. At node $B$, node $A$ requests the address table snapshot of the next hop - node $C$. When node $A$ proceeds to node $C$, it requests the current address table of node $C$ and compare it with the snapshot $A$ receives from $B$. If the difference is higher than a threshold $\tau$, $A$ may issue a fraud proof against $C$. The threshold $\tau$ is predefined by the system. Since nodes are required to update their address table snapshots with their neighbors promptly, node $C$ cannot refute the VRW fraud proof.
4.3. Table Consistency Checks
A dishonest node can store multiple copies of address tables, and use different address tables to handle different requests from different nodes for various purposes (e.g., traffic attraction, adversarial routing, etc.). We refer to the situation where a node uses more than one address table as ”equivocal tables.” To prevent equivocal tables, we employ table consistency checks (TCC). A node $A$’s random walk message carries $A$’s address table and encounter table. For each hop of the random walk, the node $B$ at that hop can compare its address table and encounter table with those of node $A$. If there is an overlapping node $C$ in $A$’s tables and $B$’s tables, $A$ and $B$ will further examine the address table snapshots of node $C$ to check if there is any difference. If the difference is higher than a threshold $\tau_{t}$, $A$ and $B$ will issue a fraud proof against node $C$ and the system may choose to slash node $C$. Similar to VRW, the threshold $\tau_{t}$ is predefined by the system. If the ratio of different peers in two address table snapshots of node $C$ within time $t$ exceeds $\tau_{t}$, then we consider the evidence is significant enough for nodes to issue a fraud proof for node $C$. Node $C$ can refute the TCC fraud proof against it only if it can provide legitimate random walk history that can explain the difference.
5. Evaluation
In this section, we evaluate Honeybee  and the baseline algorithms and compare their performances. In §5.1, we introduce the baseline algorithms and present the experimental setup. We then evaluate Honeybee  and the baseline algorithms from different perspectives and provide the results in §5.2.
5.1. Experimental Setup
We describe the baseline algorithms in §5.1.1, network layouts in §5.1.2, adversary configurations §5.1.3, and adversary strategies in §5.1.4.
5.1.1. Baselines
Since the main goal of Honeybee  is to conduct near-uniform peer sampling in p2p networks under the data availability setting, we compare Honeybee  with two arguably most potential candidate algorithms under such setting (Neu, 2022) - Kademlia (Maymounkov and Mazieres, 2002) and GossipSub (Vyzovitis et al., 2020).
  • Kademlia: Kademlia is one of the most popular p2p protocol in today’s Internet. It is used in various systems including Ethereum, Swarm, Storj, IPFS, etc. A Kademlia node has a binary node ID that is randomly assigned when the node joins the network. To route messages, each Kademlia node has a routing table consists of ”k-buckets”, and the number of k-buckets it has is equal to the length of its node ID. The $i$-th k-bucket of node A stores peers with node IDs that share the first $i-1$ bits with node A. A Kademlia node discovers new peers mainly with lookups. When node $A$ performs a lookup on node ID $B$, $A$ sends the lookup message to the neighbor(s) whose node ID is closest to $B$ in terms of XOR distance. The neighbor(s) and the nodes on the subsequent hops repeat this procedure until they find the closest node(s) to $B$. For more details about Kademlia, we refer the reader to the Kademlia paper (Maymounkov and Mazieres, 2002).
  • GossipSub: GossipSub is arguably the most renowned publish-subscribe gossip network in today’s Internet. It is used in libp2p, an open source project from IPFS. A GossipSub node has its mesh connections and gossip connections. The mesh connections are bidirectional connections, and nodes use them to send full messages. The gossip connections are unidirectional, and nodes use them to send metadata only. A GossipSub node discovers new peers mainly with peer exchanges, in which a node shares the information of some of the peers it knows with others. For more details about GossipSub, we refer the reader to the GossipSub paper (Vyzovitis et al., 2020).
5.1.2. Network
We built a discrete-event network simulator using Python based on the model description in §3.1. We simulate three p2p networks - the Kademlia network, the GossipSub network, and the Honeybee  network. Each of these networks consists of 16,384 ($=2^{14}$) nodes. The Kademlia network simulates the vanilla Kademlia network as discussed in §5.1.1. The Kademlia nodes each has 14 k-buckets and each k-bucket has size of 3 (note that some buckets may never have enough peers to reach the size limit), and we use $\alpha=3$ for all lookups. The GossipSub network simulates one overlay of the vanilla GossipSub network as discussed in §5.1.1. We simulate peer discovery in GossipSub with peer exchanges. We also simplified the scoring function by giving every peer the same score to fit GossipSub into the data availability sampling background. The scoring function from GossipSub may not help honest nodes in the data availability sampling setting. The scoring function uses parameters such as time in mesh, first message deliveries, and mesh message delivery rates. With these parameters, dishonest nodes can strategically exploit the scoring function and behave very well in terms of scores before eclipsing an honest node. The GossipSub nodes use $D\_high=12$, $D=8$, and $D\_low=6$, which are the same as the values used in the vanilla GossipSub paper. The Honeybee  network simulates Honeybee  as discussed in §4. For the Honeybee  network, 17 nodes ($\approx 1\%$) are the bootstrap nodes. Similarly, the Honeybee  nodes each uses an address table size of 24. The address table of a Honeybee  node consists of the outgoing address table and the incoming address table, each contains at most 12 nodes. In all three networks, the routing table of a node is defined as all the peers whom it knows and has access to (i.e., stores their IP addresses), and we assume there is no churn from the joining and leaving of nodes.
5.1.3. Adversary
For each of the three p2p networks, we simulate scenarios in which attackers control 5%, 10%, 20%, 30%, 40%, and 50% of the total nodes. The attacker-controlled dishonest nodes (Sybil nodes) are randomly chosen from all the nodes at the beginning of the simulation. To disturb the network and compromise potential victim node(s), the Sybil nodes employ strategies addressed in §5.1.4 during the simulation. Attackers have full control of their Sybil nodes and are able to build a desired topology with their Sybil nodes. We consider three types of initial Sybil node layouts as follows.
  • Mixed layout: the Sybil nodes are mixed into the p2p network. All the nodes in the network are connected to each other in a random way, and the Sybil nodes are simply among them. We consider this layout the most natural initial attacker layout.
  • Big cluster layout: the Sybil nodes form a big cluster. Most (98%) of the Sybil nodes only connect to each other while 2% of the Sybil nodes have connections to the honest cluster. The Sybil nodes whose connections include a connection to the honest nodes are named gateway Sybil nodes. The Sybil nodes who do not have a connection with the honest nodes are named trap Sybil nodes. The two types of Sybil nodes may switch their roles during the simulation.
  • Small clusters layout: the Sybil nodes form multiple small clusters. The Sybil nodes form 100 clusters each consists of roughly 81 nodes. Each Sybil cluster maintains one connection with the honest cluster. In other words, each Sybil cluster has one gateway Sybil nodes and the remaining nodes in the cluster are trap Sybil nodes.
We consider two types of victims: single victim node and multiple victim nodes. In the first case, Sybil nodes target a random honest node throughout the simulation (e.g., attackers attempt to compromise a client). In the second case, Sybil nodes target all honest nodes throughout the simulation (e.g., attackers attempt to compromise an organization or a company). In both cases, the goal of attackers is to eclipse the victim(s) by inserting as many Sybil nodes into the routing tables of the victim(s) as possible.
5.1.4. Adversary Strategies
Attackers can employ a wide range of adversary strategies using the nodes they control. Listed below are the strategies that we consider to be the most important. We categorize the strategies into two types: active strategies and passive strategies. We define the following active adversary strategies.
  • Request flood: attackers can repeatedly send requests (e.g., connection requests) to the victim node(s) from their pseudonymous identities until the victim node(s) get eclipsed or the attackers achieve their goal through other means. It is difficult for the honest nodes to detect and defend against this strategy since, in a permissionless p2p network, attackers may own a large number of pseudonymous identities and can send requests from different identities. It is challenging to distinguish the dishonest nodes from the honest nodes before attackers cause actual damage to the victim node(s).
  • Adversary routing: once a request reaches a dishonest node, the dishonest node may route the request in its favor to achieve certain goals. The goals include but are not limited to: preventing the request from reaching its intended destination, guiding the request to a dishonest destination, causing overhead (e.g., delays) for the request initiator, and causing inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator.
  • Adversary peer selection: an honest node should select its peers in a random way or based on certain bona fide rules. An honest node should not make its peer selection decisions based on the identity of the candidate peers (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the candidate peers. For example, a group of dishonest nodes may choose to add each other to their routing tables to form a Sybil cluster.
  • Equivocal table: a dishonest node can keep multiple copies of routing tables, each storing different peers, and use different routing tables to route different requests. This strategy can prevent the request from reaching its intended destination, guide the request to a dishonest destination, help other dishonest nodes with load-balancing, cause overhead (e.g., delays) for the request initiator, and cause inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator.
We define the following passive adversary strategies.
  • Selective request acceptance: an honest node should accept/reject a connection request in a random way or based on certain bona fide rules. An honest node should not make its acceptance/rejection decisions based on the identity of the request initiator (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the request initiator. For example, attackers and their pseudonymous identities may choose to reject requests from all the honest nodes except for its targeted victim node(s).
  • Adversary recommendation: upon request, an honest node should recommend (i.e., share the information of) a peer to another node in a random way or based on certain bona fide rules. An honest node should not make its recommendation decisions based on the identity of the request initiator and (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the request initiator. For example, an honest victim’s dishonest neighbor may choose to recommend other dishonest nodes to the victim. This strategy is similar to request flood but in a passive manner.
  • Black hole: once a request reaches a dishonest node, the dishonest node may drop the request completely (i.e., being unresponsive to the request). This strategy can cause overhead (e.g., delays) and inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator. In comparison with adversary routing, this strategy is less harmful but more difficult to detect/deter since honest nodes can also be unresponsive for various legitimate reasons.
5.2. Results
(a) Honeybee
(b) GossipSub
(c) Kademlia
Figure 2. Nodes adhere to protocol: Sampling distribution from a random observation node in 100 thousand epochs. Every node follows its protocol. Node IDs are sorted by frequency in ascending order. True uniform sampling distribution is shown as a dashed line.
5.2.1. Sampling Distributions
Fig. 2a, 2b, and 2c show the sampling distributions for an arbitrarily chosen observation node with Honeybee , GossipSub, and Kademlia in a network of 16,384 honest nodes for 100 thousand epochs. In each setting, all nodes start (i.e., epoch 0) with random routing table configurations, and all nodes behave according to their protocol. In the figures, the true uniform sampling distribution (i.e., all nodes except for the observation node are sampled with equal probability) is shown as a dashed line. We observe that Honeybee  and the two baseline algorithms have similar sampling distributions when all nodes adhere to their protocol.
(a) Total variation distances
(b) Chi-Square test results
(c) Chi-Square with 30% idle nodes
Figure 3. Quality of sampling when nodes adhere to protocol: (a) displays comparisons of Honeybee, GossipSub, and Kademlia with the true uniform sampling distribution; (b) displays the Chi-Square test results with p-value of 0.05 shown as a dashed line; (c) displays the Chi-Square test results when 30% of the nodes are idle.
To compare the sampling distributions of Honeybee  and the baseline algorithms with the true uniform sampling distribution, Fig. 3a plots the total variation distances between the three algorithms and the true uniform sampling distribution from the above experiment. The total variation distance curves suggest that sampling distributions from Honeybee  and the baseline algorithms converge to the true uniform sampling distribution throughout the 100 thousand epochs in a similar pattern. At the beginning of the simulations, the total variation distances from Honeybee  and the baseline algorithms are higher than 95%. At 100,000 epochs, the total variance distances from Honeybee  and the baseline algorithms decrease to roughly 23%. We also extended the simulations to 3 million epochs, and the total variation distances from the three algorithms decrease to 3-5%.
To examine the sampling distributions in detail, we conduct Chi-Square tests on the samples from Honeybee  and the baseline algorithms. We combine neighboring cells by dividing the node IDs into 127 node ID intervals of equal length (since the number of node IDs minus the observation node is divisible by 127). We divide the 100 thousand epochs into 10 time intervals of equal length. For each time interval, we examine the samples from that particular time interval. In Fig. 3b, we plot the Chi-Square values across the 10 time intervals for Honeybee  and the baseline algorithms. In Fig. 3c, we plot the Chi-Square values across the 10 time intervals for Honeybee  and the baseline algorithms when 30% of the nodes are idle (i.e., do not actively conduct sampling). We observe that, for Honeybee  and the baseline algorithms, we cannot reject the null hypothesis of uniform sampling with sufficient evidence when all nodes conduct sampling. When there are 30% of idle nodes, the Chi-Square values slightly increase for Honeybee  and GossipSub.
5.2.2. Why VRW and TCC
We demonstrate the importance of verifiable random walks and table consistency checks by showing the reader what happens when we do not use verifiable random walks or table consistency checks in Honeybee. Without the random walk verification mechanism, a random walk is not necessary random since dishonest nodes can hijack it by reconfigure their address tables according to their preferences. Similarly, without the table consistency check mechanism, dishonest nodes are able to store multiple copies of address tables and use different copies to route random walks according to their preferences. In other words, without verifiable random walks or table consistency checks, dishonest nodes can carry out a de facto adversarial routing attack.
(a) Honeybee  without VRW
(b) Honeybee  without TCC
Figure 4. Honeybee  without VRW or TCC for one random victim under attack: (a) displays comparison of vanilla Honeybee  and Honeybee  without random walk verification mechanism; (b) displays comparison of vanilla Honeybee  and Honeybee  without table consistency check mechanism.
We simulate two altered versions of Honeybee: one without the random walk verification mechanism, the other without the table consistency check mechanism. Fig. 4a shows the process of dishonest nodes target and eclipse a random honest victim node when we remove the verification mechanism from random walks. Fig. 4b shows the process of dishonest nodes target and eclipse a random honest victim node when we remove table consistency checks. In both scenarios, all nodes start with a random address table configuration, and we have 30% of dishonest nodes in the network. We compare them with the unaltered Honeybee  under the same setting to present the consequences.
(a) 5% dishonest nodes
(b) 10% dishonest nodes
(c) 20% dishonest nodes
Figure 5. An honest node attacked by 5%, 10%, and 20% of dishonest nodes: Single random honest node under attack in Honeybee, GossipSub, and Kademlia.
(a) 30% dishonest nodes
(b) 40% dishonest nodes
(c) 50% dishonest nodes
Figure 6. An honest node attacked by 30%, 40%, and 50% of dishonest nodes: Single random honest node under attack in Honeybee, GossipSub, and Kademlia.
5.2.3. Attack - Single Victim
Fig. 5a, 5b, and 5c show the ratio of dishonest peers in a random honest victim node’s routing table in 1,000 epochs for Honeybee  and the baseline algorithms in a network of 16,384 nodes, and 5%, 10%, and 20% of the total population consists of dishonest nodes. Fig. 6a, 6b, and 6c show the ratio of dishonest peers in a random honest victim node’s routing table in 1,000 epochs for Honeybee  and the baseline algorithms in a network of 16,384 nodes, and 30%, 40%, and 50% of the total population consists of dishonest nodes. Each node starts with a random routing table configurations. We observe that, under different levels of dishonest node percentage, Honeybee  outperforms the baseline algorithms in terms of the mean ratio of dishonest peers in the victim’s routing table.
Percent of dishonest nodes in the network
5% 10% 20% 30% 40% 50%
Algorithm Honeybee 6.16% 11.49% 21.91% 32.25% 40.90% 51.14%
GossipSub 9.99% 25.12% 42.24 55.52% 69.64% 81.79%
Kademlia 56.26% 63.22% 85.12% 87.15% 97.36% 95.23%
Table 1. Mean ratio of dishonest peers in victim’s routing table: Five honest nodes are randomly sampled as victims in five separate simulations with Honeybee, GossipSub, and Kademlia. They are attacked by 5%, 10%, 20%, 30%, 40%, and 50% of dishonest nodes.
To demonstrate that the results apply across different nodes in different simulations, we randomly sample five honest nodes as victims in five separate simulations. We calculate the mean ratio of dishonest peers in the victim’s routing table among the five nodes over the course of 1,000 epochs and present the results in Table. 1. We observe that Honeybee  consistently achieves near-uniform sampling and outperforms the baseline algorithms by 4-63% in terms of the mean ratio of dishonest peers in the victim’s routing table.
Figure 7. All honest nodes under attack: In a network of 16,384 nodes, 50% nodes are dishonest nodes. The honest nodes run Honeybee, GossipSub, and Kademlia. The dishonest nodes attempt to eclipse as many honest nodes as possible over the course of 1,000 epochs.
5.2.4. Attack - Multiple Victims
Fig. 7 shows the cumulative number of honest nodes that can be eclipsed by dishonest nodes in a network with 50% dishonest nodes in 1,000 epochs with Honeybee, GossipSub, and Kademlia. There are 8,192 honest nodes in total. We observe that, in 1,000 epochs, dishonest nodes can eclipse over 75% of honest nodes with GossipSub and Kademlia while they cannot eclipse honest nodes with Honeybee.
6. Related Work
6.1. p2p Networks
Whether the purpose is for enhancing blockchain data availability sampling or not, researchers have made extensive effort to improve the security and efficiency of p2p networks. In Castro et al. (Castro et al., 2002), the authors proposed secure routing against attacks that target p2p network message deliveries. The secure routing is achieved with restricted node ID assignment, constrained routing table, and diverse delivery routes. In Baumgart and Mies (Baumgart and Mies, 2007), the authors presented a secure Kademlia key-based routing protocol. They limited free node ID generation with a supervised signature or a crypto puzzle signature and introduced a reliable sibling broadcast. In Coretti et al. (Coretti et al., 2022), the authors designed a byzantine-resilient gossip protocol under the proof-of-stake setting. Under the protocol, nodes build a connected backbone of high-staking nodes. However, light nodes (or nodes with small stakes) are easily sacrificed in this protocol, and stake distribution is not always readily available in a completely decentralized environment. Recent researches inspired by the multi-armed bandit problem (Mao et al., 2020; Wei and Bojja Venkatakrishnan, 2023; Zhang and Bojja Venkatakrishnan, 2023) attempted to reduce the communication latency in p2p networks without losing security by learning from the performance of neighboring peers and update peer selections based on the knowledge. In Kiffer et al. (Kiffer et al., 2021), the authors studied the connectivity and block propagation mechanism of Ethereum’s p2p network. In Król et al. (Król et al., 2023), the authors conducted an analysis on p2p networking requirements for data availability sampling in Ethereum.
6.2. Data Availability Sampling
Plenty of efforts have been made in blockchain layer-2 scaling (Ethereum.org, 2024b, e; Sivaraman et al., 2020; Zhang and Venkatakrishnan, 2023). As a critical component of layer-2 solutions, data availability sampling itself has many important research problems (Neu, 2022). In Hall-Andersen et al. (Hall-Andersen et al., 2023), the authors initiated a cryptographic study of data availability sampling and demonstrated its relation with erasure codes. In Al-Bassam et al. (Al-Bassam et al., 2021), the authors developed a complete fraud and data availability proof scheme in which they claimed that light nodes can have a security guarantee close to the level of a full node under certain assumptions. In Yu et al. (Yu et al., 2020), the authors created coded Merkle trees to improve the protection of light nodes against data availability attacks. In Cao et al. (Cao et al., 2020), the authors presented a decentralized collaborative light-node-only verification mechanism in which light nodes can conduct block verification without the help of a full node. In addition to the advancement in security, there are some researches that focus on lowering the communication and computation overhead on nodes in the data availability setting (Sheng et al., 2021; Kiayias et al., 2020; Bünz et al., 2020; Kiayias et al., 2016).
7. Conclusion
We presented Honeybee, a decentralized p2p sampling algorithm that achieves near-uniform peer sampling for blockchain data sharding. Different from existing algorithms, Honeybee nodes conduct genuine random walks in the network using verifiable random walks and table consistency checks, effectively filled the void on the p2p layer for data availability sampling. Honeybee is secure against various adversarial strategies under diverse settings. In our experiments, we observe that Honeybee consistently achieves $\epsilon$-uniform sampling with $\epsilon=0.03$. An interesting future research topic would be to generate a Markov chain using Honeybee nodes and their address tables. Then, we could examine from a theoretical perspective whether mixing occurs and the potential mixing time when all nodes actively perform random walks.